OpenStudy (anonymous):

Identify the 25th term of the arithmetic sequence 2, 1 and 3 over 5, 1 and 1 over 5 … @ranga

OpenStudy (kc_kennylau):

1. Find the common difference d. 2. Calculate the value of a+24d (a is the first term)

OpenStudy (kc_kennylau):

rmb each arithmetic sequence has a common difference :) the first term is a the second term is a+d the third term is a+2d the nth term is a+(n-1)d

OpenStudy (anonymous):

is the common difference is -2/5

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

yes

OpenStudy (kc_kennylau):

then what's the 25th term?

OpenStudy (anonymous):

Ok hold up.

OpenStudy (kc_kennylau):

ok

OpenStudy (anonymous):

-19 1/5

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

um no

OpenStudy (anonymous):

umm lol what did i do wrong

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

How did you get your answer?

OpenStudy (anonymous):

never mind let me try agaain i saw my mistake

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

nope, you forgot to minus one (a+24d not a+25d)

OpenStudy (anonymous):

ugh so what would it be

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (anonymous):

i did it with 24 idk what ur talking about @kc_kennylau

OpenStudy (kc_kennylau):

Can you show me your steps? :)

OpenStudy (anonymous):

24*-2/5= -9 3/5+ 1 3/5=-8

OpenStudy (anonymous):

@kc_kennylau can you reply with not having to tag you each time

OpenStudy (kc_kennylau):

sorry i'm doing two questions at once, so please tag me

OpenStudy (kc_kennylau):

How did you get -9 3/5 and 1 3/5 why are you complicating everything

OpenStudy (anonymous):

pls can you jst show me ur steps @kc_kennylau

OpenStudy (kc_kennylau):

Common difference =\(1\dfrac35-2\) =\(\dfrac25\) The 25th term =\(a+24d\) =\(2-24\times\dfrac25\) =\(2-\dfrac{48}5\) =\(2-9\dfrac35\) =\(-\left(9\dfrac35-2\right)\) =\(-7\dfrac35\)

OpenStudy (anonymous):

so if you dont follow order of operation you would get this wrong? @kc_kennylau

OpenStudy (kc_kennylau):

oh i see what you did there

OpenStudy (kc_kennylau):

you used the second term to add the 24*difference

OpenStudy (kc_kennylau):

you should use the first term

OpenStudy (anonymous):

thank you